You measure 21 textbooks' weights, and find they have a mean weight of 43 ounces. Assume the population standard deviation is 12.6 ounces. Based on this, construct a 90% confidence interval for the true population mean textbook weight.

Give your answers as decimals, to two places

< μ <
Answer #1

Solution :


Given that,

Point estimate = sample mean = = 43

Population standard deviation =    = 12.6

Sample size = n =21

At 90% confidence level

= 1 - 90%  

= 1 - 0.90 =0.10

/2 = 0.05

Z/2 = Z0.05 = 1.645( Using z table )   


Margin of error = E = Z/2 * ( /n)

= 1.96* ( 12.6 /  21 )

E= 4.52
At 90% confidence interval estimate of the population mean
is,

- E < < + E

43 - 4.52 <   < 43+ 4.52

38.48 <   < 47.52